OpenStudy (anonymous):

Suppose that x and y vary inversely and that x = 2 when y = 8. Write the function that models the inverse variation

OpenStudy (anonymous):

y = 16/x y = 10/x y = 4x y = 6/x

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):



OpenStudy (anonymous):

when y=8 , x=2 k/8=2 k=16

OpenStudy (anonymous):

So, x=16/y

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

k got it thank youu!! <333333 wait but isn't it y = 16/x

OpenStudy (anonymous):

same thing